Solve the following system of equations.
x+2y=7 (1)
x=8-2y (2)
How do I do this one???

Answer by rapaljer(4671)

Since the second equation is stated in the form "x= ____", I would recommend the substitution method. Substitute what x= in the second equation, back into the first equation:

x+2y=7 (1)
x=8-2y (2)

(_____) +2y = 7
(8-2y) +2y=7

A funny thing happens here!! All the variables subtract out, and you are left with 8=7. This is NOT POSSIBLE, so there is NO SOLUTION. Actually, the two given straight lines are PARALLEL LINES.

For a more detailed explanation, see my own website by clicking on my tutor name "rapaljer" anywhere in algebra.com. Click on the second link on my homepage, which is "Basic, Intermediate, and College Algebra: One Step at a Time." Click on Basic Algebra, and choose Chapter 4. Problems like this one will be found in Section 4.07 Systems of Equations. You may find my explanations easier to understand than traditional textbooks.

R^2